OpenStudy (deadshot):

How do I find the 16th term of a geometric sequence where a{subscript(1)} = 4 and a{subscript(8)} = -8,748.

OpenStudy (deadshot):

\[a _{1}=4\]and\[a _{8}=-8,748\]

OpenStudy (amistre64):

since you know a1, and the general formula for a geometric sequence is:\[a_n=a_1~r^{n-1}\] you can determine r readily

OpenStudy (anonymous):

\[4\times r ^{7}= -8748\]

OpenStudy (anonymous):

\[\sqrt[7]{-2187} = -3 = r \]

OpenStudy (anonymous):

\[ar ^{15}= 4 \times (-3)^{15}= -57395628\]

OpenStudy (deadshot):

Thanks!